OpenStudy (anonymous):

find the inverse of each function. then graph the function and its inverse. g(x)=2x^2-3

OpenStudy (luigi0210):

Switch x and y, then solve for y again. So \(\Large y=2x^2-3 ======> x=2y^2-3\)

OpenStudy (anonymous):

how do you graph it?

OpenStudy (luigi0210):

Do you have the inverse equation?

OpenStudy (anonymous):

is it y=i sq root 2x+6/2?

OpenStudy (luigi0210):

I don't think you need an i there. \(\Large f^{-1}(x)=\sqrt{\frac{x+3}{2}}\)

OpenStudy (anonymous):

so how'd u graph it?

OpenStudy (luigi0210):

You can just plot in random x-coordinates and get out a y-coordinate. Here's what it would look like:

OpenStudy (anonymous):

what about if its a fraction--> -2/3x

OpenStudy (luigi0210):

\(\Large -\frac{2}{3x}\) or \(\Large -\frac{2}{3}x\)?

OpenStudy (anonymous):

the second one

OpenStudy (luigi0210):

It's the same as \(\Large -\frac{2x}{3}\) Just multiply. So say you had \(x=3\) you plug that in \(\Large -\frac{2}{3}*3\) The 3 cancels, and you are left with -2. Or if you had 2. \(\Large -\frac{2}{3}*2\) It would be \(\Large -\frac{4}{3}\).

OpenStudy (anonymous):

what would be the inverse of the equation pls?

OpenStudy (anonymous):

Here are some steps that'll help you, To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (luigi0210):

Oh, inverse? \(\Large y=-\frac{2x}{3}===> x=-\frac{2y}{3}\) Solve for y by multiplying by 3, the dividing: \(\Large -\frac{3x}{2}=f^{-1}(x)\)